Find the inverse of g(x)=-2x+4
Answer by Cromlix(4381)

Hi there,
g(x) = -2x + 4
y = -2x + 4
Change the y and the x around
x = -2y + 4
x - 4 = -2y
(x - 4)/-2 = y
y = (x - 4)/-2
This is the inverse.
Hope this helps :-)